The mathematical scores of 6 groups of 4 students each are shown below. Test whether the differences in scores may be attributed to chance, at .01 significance level.

 Group 1: Student1 84, Student2 88, Student3 70, Student4 80. Group 2: Student1 90, Student2 95, Student3 93, Student4 80. Group 3: Student1 70, Student2 73, Student3 85, Student4 90. Group 4: Student1 95, Student2 96, Student3 90, Student4 90. Group 6: Student1 85, Student2 78, Student3 75, Student4 90. Group 6: Student1 93, Student2 85, Student3 80, Student4 90.


1
Expert's answer
2021-11-16T17:38:27-0500

One-Way ANOVA test

using online calculator https://goodcalculators.com/, we get:


H0: µ1 = µ2 = µ3    ("all 3 population means are equal")

H1: At least one µi different ("at least one of the 3 population means is not equal to the others")

test statistic:

computed as

"F=\frac{MSR}{MSE}"

where MSR = SSR/dfr = the regression mean square,

MSE = SSE/dfe = the mean square error,

SSR = the regression sum of squares,

SSE = the error sum of squares,

SST = the total sum of squares (SST = SSR + SSE),

dfr = the model degrees of freedom (equal to dfr = k - 1),

dfe = the error degrees of freedom (equal to dfe = n - k - 1)

k = the total number of groups (levels of the independent variable)

n = the total number of valid observations,

dfT = the total degrees of freedom (equal to dfT = dfr + dfe = n - 1)


p-value > 0.01

so, we can conclude that the differences in scores may be attributed to chance